# The immersed boundary method: an accelerated SIMPLE approach for moving bodies

### v1

@@ 2.2 The SIMPLE approach for IBM
First, the intermediate velocity field ${\bf u}^{*}$ is obtained from the solution of Eq. (8). We note that this velocity field does not necessarily satisfy the no-slip and divergence-free constraints. Second, a divergence is applied on both sides of Eq. (9), similarly to the original SIMPLE method, yielding the modified pressure Poisson equation
$-\nabla^{2}p+\nabla\cdot(R{\bf F}^{\prime})=-\frac{3}{2\Delta t}\nabla\cdot{\bf u}^{*},$ (12)
@@
where the right hand side in this equation is achieved by utilizing the divergence free constraint in Eq. (11), that implies $\nabla\cdot{\bf u}^{*}=-\nabla\cdot{\bf u}^{\prime}$. However, Eq. (12) for the pressure corrections is not closed, as it also includes the forces corrections field. Thus, utilizing again Eq. (9) we obtain
@@
By applying the interpolation $R^{T}$ on both sides of Eq. (13) and substituting it back into Eq. (10), we attain the following system of equations governing the corrections for the pressure and forces fields:
@@
where $G$ is the discretized gradient, $RHS_{p^{\prime}}=-\frac{3}{2\Delta t}\nabla\cdot{\bf u}^{*}$ and $RHS_{{\bf F}^{\prime}}=\frac{3}{2}\frac{R^{T}{\bf u}^{*}-{\bf U}^{\Gamma}}{\Delta t}$. This system was solved in [16] by the block reduction
$S_{d}{\bf F}^{\prime}$ $=BL^{-1}RHS_{p^{\prime}}-RHS_{{\bf F}^{\prime}}$ (15)
$p^{\prime}$ $=L^{-1}(RHS_{p^{\prime}}-B^{T}{\bf F}^{\prime}),$ (16)
where $S_{d}$ is the dual Schur-complement:
$S_{d}\coloneqq C+BL^{-1}B^{T}.$ (17)
Their approach yielded the ${\bf F}^{\prime}$ field by solving Eq. (15) via either direct [25] or Krylov iteration [26] solvers, followed by calculation of the $p^{\prime}$ field by Eq. (16). While $S_{d}$ is relatively small, constructing it for realistic 3D configurations requires applying the Laplacian solver hundreds of thousands of times (for each column of $B^{T}$, corresponding to a Lagrangian point). Although this task is feasible as a precomputing step for stationary immersed bodies, it becomes impractical for moving boundary simulations where $B^{T}$ changes at each time step.
3 Method
@@ 3.1 The Laplacian as a preconditioner
To accelerate the suggested method and adjust it for moving boundary simulations, we propose the following remedy: instead of forming the dual Schur complement, we form the primal Schur complement
@@
and solve the resulting decoupled system of equations, first for the pressure $p^{\prime}$ and then for the force $\textbf{F}^{\prime}$ corrections fields:
$S_{p}p^{\prime}$ $=RHS_{p^{\prime}}+B^{T}C^{-1}RHS_{{\bf F}^{\prime}}$ (19)
${\bf F}^{\prime}$ $=C^{-1}(Bp^{\prime}-RHS_{{\bf F}^{\prime}}).$ (20)
Forming the primal Schur complement $S_{p}$ and calculating the right hand side in Eq. (19) and (20) is significantly cheaper than forming the dual Schur complement $S_{d}$ and calculating the right hand side in Eq. (15) and (16): in the virtue of the scalar approximation $C\approx\frac{1}{2}I$, that was suggested in [16], we avoid calculating $C^{-1}$ and use the approximation $C^{-1}\approx 2I$ and
@@
This approximation is valid for immersed bodies with sufficiently smooth surfaces and improves with increasing grid resolution. After applying this scalar approximation on both Eq. (19) and (20), we obtain
$\tilde{S}_{p}p^{\prime}$ $=RHS_{p^{\prime}}+2B^{T}RHS_{{\bf F}^{\prime}}$ (22)
${\bf F}^{\prime}$ $=2(Bp^{\prime}-RHS_{{\bf F}^{\prime}}).$ (23)
However, both $S_{p}$ and $\tilde{S}_{p}$ are ill-conditioned and large, and hence neither a direct nor an iterative solution of Eq. (19) (or the approximated version, Eq. (22)) is practical, and a good preconditioner is required. To this end, we use the Laplacian $L$ as a preconditioner. We show in the following sections that Eq. (19) can be solved in just a few iterations of a preconditioned Krylov method, because $L$ and $S_{p}$ are spectrally equivalent. That is, we prove in Section 4 that the eigenvalues of the preconditioned operator $L^{-1}S_{p}$ lie in a narrow constant interval, that does not depend neither on the physical variables nor on the grid size. As long as $\tilde{S}_{p}$ well approximates $S_{p}$, this result nearly holds for the preconditioned operator $L^{-1}\tilde{S}_{p}$, and as we show numerically in Section 5, Eq. (22) can be solved in just a few iterations. Since the solution of the pressure-forces corrections system comprises the computational bottleneck, our preconditioning approach enables a fast solution of the problem as a whole, and favorable scalability properties.
The resulting method is summarized in Alg. 1.
  1. 1.
    Solve Eq. (8) for ${\bf u}^{*}$.
  2. 2.
    Solve Eq. (22) iteratively, using a Krylov iteration based method preconditioned by $L$, for $p^{\prime}$.
  3. 3.
    Substitute the obtained $p^{\prime}$ in Eq. (23), and calculate $\textbf{F}^{\prime}$.
  4. 4.
    Complete the pressure correction $p^{(n+1)}=p^{(n)}+p^{\prime}$.
  5. 5.
    Compute ${\bf u}^{\prime}$ by substituting the obtained $p^{\prime}$ and $F^{\prime}$ into Eq. (9).
  6. 6.
    Complete the velocity projection step ${\bf u}^{(n+1)}={\bf u}^{*}+{\bf u}^{\prime}$.
Algorithm 1 accelerated SIMPLE for IBM.
@@
There is abundance of literature on block-preconditioning for saddle point matrices as appears in Eq. (14), mainly by Schur complement preconditioning. Since forming and inverting the exact Schur-complement is computationally expensive, broad research is dedicated to the search of cheap approximations for the inverse of the Schur-complement, see [27]. Our method offers an efficient approximation for the primal Schur complement for a certain class of saddle point matrices, and can be implemented either by Alg. 1 or by a corresponding block lower triangular Schur complement preconditioner.
3.2 Implementation details
In this subsection we describe our efficient matrix-free implementation. First, the operator $B^{T}$ consists of the (minus) divergence applied to the discrete Dirac delta function’s weight terms obtained by Eq. (5). A matrix free implementation allows efficient shared-memory parallelization (we used OpenMP) and enables rapid re-computation of these operators at each time step. Second, for each step of the preconditioned Krylov iteration, only a matrix-vector product against the preconditioned operator is needed. We use left preconditioning, which is mathematically equivalent to solving Eq. (22) with both sides multiplied by $L^{-1}$ from the left. However, we don’t need to form $L^{-1}$, nor to apply a Laplacian solver at multiple columns, in order to perform a preconditioning step. Since the matrix of the preconditioned system is
@@ 3.2 Implementation details
where $I_{m}$ is an identity matrix of the appropriate size, the product of $L^{-1}\tilde{S}_{p}$ with a vector ${\bf v}$ can be conveniently obtained with the direct solver suggested in [19] applied, only once, on the vector $B^{T}(B{\bf v}).$ Employing Intel MKL routines [^2] We convert $B$ and $B^{T}$ to the CSR format at each time step for further utilization of Intel MKL routines. enables efficient calculation of the sparse matrix-vector products. Note that since the scalar approximation for $C^{-1}$ is used, the calculation of the second term in Eq. (22) and of the entire Eq. (23) are trivial, as they both require a single matrix-vector product.
Finally, we discuss the factorization of the preconditioner $L$. It was suggested in [19] that any matrix that can be written as a sum of Kroncker products of smaller matrices with identity matrices can be factorized by utilizing an eigen-decomposition of each of the small matrices. As shown there, the system can be solved with a complexity of $O(n^{4/3})$ for a regular 3D grid. The computational cost can be further reduced by utilizing the Thomas TDMA algorithm in one of the three directions. For transient problems this method is highly efficient, as shown in [20], because the corresponding operators can be factorized in advance and used in each time step.

### v2

@@ 3.1 One-way coupled moving boundary
When proceeding with the solution, the intermediate velocity field $\bf u$$^{*}$ is first obtained from the solution of Eq. (14). We note that this velocity field does not necessarily satisfy the no-slip and divergence-free constraints. Second, a divergence is applied to both sides of Eq. (15), similarly to the original SIMPLE method [41], yielding the modified pressure Poisson equation:
$\nabla^{2}p+\nabla\cdot(R{\bf F}^{\prime})=-\frac{3}{2\Delta t}\nabla\cdot{\bf u}^{*},$ (18)
@@
where the right-hand side is obtained by utilizing the divergence-free constraint in Eq. (17), which implies $\nabla\cdot{\bf u}^{*}=-\nabla\cdot{\bf u}^{\prime}$. However, Eq. (18) for the pressure correction is not closed, as it also includes the force correction field. Thus, using Eq. (15) again, we obtain:
@@
By applying the interpolation operator $R^{T}$ to both sides of Eq. (19) and substituting the result into Eq. (16), we obtain the following system of equations governing the corrections for the pressure and force density fields:
@@
where $G$ is the discretized gradient operator, $RHS_{p^{\prime}}=-\frac{3}{2\Delta t}\nabla\cdot{\bf u}^{*}$, and $RHS_{{\bf F}^{\prime}}=\frac{3}{2\Delta t}(R^{T}{\bf u}^{*}-{\bf U}^{\Gamma})$. Once the system in Eq. (20) is solved, we update the velocity, pressure, and Lagrangian force density fields and proceed to the next time step. [^3] In principle, internal fixed-point iterations may be applied to improve accuracy. However, for one-way coupling configurations and a sufficiently small time step ($\Delta t=10^{-3}$), the desired accuracy is achieved in a single step [30]. The system of equations in Eq. (20) constitutes a large, regularized saddle-point problem, whose efficient solution is a major bottleneck of the method described for one-way moving boundary, and in the next subsection we derive a similar equation that constitutes the bottleneck for two-way FSI configurations as well. The efficient solution of this saddle-point problem is a core contribution of the present study and is addressed in detail in the following sections.
3.2 Two-way coupled FSI
The conceptual difference between one-way coupled moving boundary simulations and two-way coupled FSI lies in the fact that the latter requires coupling the NS equations, governing the fluid dynamics, with the structural equations governing the dynamics of the solid body — specifically, Eqs. (4) and (5), which represent the conservation of linear and angular momentum of the solid particle. Remarkably, both equations include contributions from fluid inertia, commonly referred to as the added mass effect, which accounts for the rate of change of linear and angular momentum of the fluid enclosed within the volume occupied by the immersed body [38]. These contributions appear as volume integrals over the domain occupied by the rigid body and must be evaluated accurately to ensure physical fidelity of the simulation. In the current study, we adopt the voxel-based strategy proposed by [40], wherein the particle surface is approximated using a Cartesian grid. In this approach, each Eulerian cell is assigned a partial volume fraction $\alpha_{i}^{(n)}\in[0,1]$ that represents the fraction of the cell lying within the particle at time $t^{(n)}$. The computation of $\alpha_{i}^{(n)}$ follows the linear reconstruction scheme introduced in [40], which uses the signed distance function to approximate the volume fraction of the particle within each cell. Accordingly, the volume integrals over the particle domain, appearing in the linear and angular momentum balance equations, are approximated as:
$\int_{V_{p}^{(n)}}\mathbf{u}^{(n)}\,dV\approx\sum_{i\in V_{p}^{(n)}}\alpha_{i}\mathbf{u}_{i}^{(n)}\,\Delta V,$ (21a)
$\int_{V_{p}^{(n)}}\mathbf{x}^{(n)}\times\mathbf{u}^{(n)}\,dV\approx\sum_{i\in V_{p}^{(n)}}\alpha_{i}\mathbf{x}_{i}^{(n)}\times\mathbf{u}_{i}^{(n)}\,\Delta V.$ (21b)
3.2.1 Iterative Coupling Algorithm
As already mentioned, our method belongs to the family of partitioned approaches; that is, the fluid and structural governing equations are solved separately and coupled through internal iterations at each time step to achieve strong fluid-structure interaction. At each time step, we employ an iterative predictor-corrector scheme. The algorithm, inspired by [23], is tailored specifically to our immersed boundary framework and proceeds as follows:
Step 1: Explicit prediction of particle position and orientation
The particle’s new position and orientation are explicitly predicted with second-order accuracy in time using a backward differentiation formula:
$\mathbf{x}^{(n+1)}=\frac{1}{3}(4\mathbf{x}^{(n)}-\mathbf{x}^{(n-1)})+\frac{2}{3}\Delta t\,\mathbf{u}_{s}^{(n)}+\mathbf{R}^{(n+1)}\mathbf{x}^{(0)}.$ (22)
These quantities remain fixed throughout the subsequent inner iterations to maintain computational efficiency.
Step 2: Eulerian velocity prediction
We solve the discretized Navier–Stokes equations once to obtain an initial prediction, $\mathbf{u}^{(k_{0})}$, for the Eulerian velocity field:
$\frac{3\mathbf{u}^{(k_{0})}-4\mathbf{u}^{(n)}+\mathbf{u}^{(n-1)}}{2\Delta t}=-\nabla p^{(n)}+\frac{1}{Re}\nabla^{2}\mathbf{u}^{(k_{0})}+R\mathbf{F}^{(n)}-(\mathbf{u}\cdot\nabla\mathbf{u})^{(n)}.$ (23)
At this stage, incompressibility and no-slip conditions are not yet enforced.
Step 3: Iterative correction procedure
An iterative correction loop indexed by $k\geq 0$ is initiated:
  • 1.
    Translational velocity update:
[…]
    $\quad+\frac{2}{\pi\theta}\left(3\int_{V_{p}^{(n+1)}}\mathbf{u}^{(k)}\,dV-4\int_{V_{p}^{(n)}}\mathbf{u}^{(n)}\,dV+\int_{V_{p}^{(n-1)}}\mathbf{u}^{(n-1)}\,dV\right)$
    $\quad-\frac{2\Delta t}{3}\cdot\frac{\theta-1}{\theta}Fr^{-2}\mathbf{e}_{z}.$ (24)
    Note that this discrete form of Eq. (4), which governs the conservation of linear momentum of the solid body, is obtained here for the spherical particle by discretizing all temporal derivatives, using a second-order backward finite difference.
  • 2.
    Rotational velocity update:
[…]
    $\quad+\frac{20}{\pi\theta}\left(3\int_{V_{p}^{(n+1)}}\mathbf{x}^{(n+1)}\times\mathbf{u}^{(k)}\,dV-4\int_{V_{p}^{(n)}}\mathbf{x}^{(n)}\times\mathbf{u}^{(n)}\,dV+\int_{V_{p}^{(n-1)}}\mathbf{x}^{(n-1)}\times\mathbf{u}^{(n-1)}\,dV\right).$ (25)
    Similarly, this discrete form of Eq. (5), which governs the conservation of angular momentum of the solid body, is obtained here for the spherical particle by discretizing all temporal derivatives, using a second-order backward finite difference.
  • 3.
    Lagrangian boundary velocity evaluation: This step is required for updating the RHS of the system in Eq. (27).
[…]
  • 4.
    Pressure and force-density correction: The coupled pressure and velocity corrections are computed at each iteration. As the iteration converges, the RHS of the system below approaches zero, and so do the correction values, since the operator is not singular.
    $\begin{bmatrix}G^{T}G&G^{T}R\\[5.0pt] R^{T}G&-R^{T}R\end{bmatrix}\begin{bmatrix}p^{\prime}\\[5.0pt] \mathbf{F}^{\prime}\end{bmatrix}=\begin{bmatrix}-\frac{3}{2\Delta t}\nabla\cdot\mathbf{u}^{(k)}\\[5.0pt] \frac{3}{2\Delta t}(R^{T}\mathbf{u}^{(k)}-\mathbf{U}^{(k)})\end{bmatrix}$. (27)
  • 5.
    Fields update and convergence check: This is essentially a Gauss–Seidel step with $\xi$ acting as an under- or over-relaxation parameter. Our numerical experiments revealed that $\xi$ can range between 0.6 for the smallest and 1.1 for the largest values of $\rho_{p}/\rho_{f}$ ratio.
    $p^{(n,k+1)}=p^{(n,k)}+\xi p^{\prime},$ (28)
    $\mathbf{F}^{(n,k+1)}=\mathbf{F}^{(n,k)}+\xi\mathbf{F}^{\prime},$
    $\mathbf{u}^{(k+1)}=\mathbf{u}^{(k)}+\xi\frac{2\Delta t}{3}\left(-\nabla p^{\prime}+R\mathbf{F}^{\prime}\right).$
    Iterations (i.e., Step3Step~3) continue until the relative $L_{1}$-norm of the incremental force-density correction falls below a specified threshold (typically $10^{-3}$), ensuring accurate enforcement of divergence-free and no-slip conditions:
    $\frac{\|\mathbf{F}^{\prime}\|_{1}}{\|\mathbf{F}^{(n,k+1)}\|_{1}}<10^{-3}.$ (29)
    Upon convergence, we proceed to the subsequent time step.
This iterative scheme ensures consistent enforcement of the incompressibility and no-slip constraints while maintaining modularity between the fluid and solid solvers. Similarly to the one-way moving boundary case, the system (27) constitutes the bottleneck of the entire method and our main contribution is the preconditioning approach to address it, presented in the next section.
4 Method: preconditioning
To motivate the need of developing an efficient preconditioner, we first explain why the main computational bottleneck in our framework is the repeated solution of the regularized saddle-point linear system coupling the pressure and force-density corrections. First, we explain the computational framework of our solver. As described in the previous section, the fluid and structural dynamics equation are solved separately, and then we either advance directly to the next time step (for one-way coupled simulations) or perform fixed-point internal iterations (for two-way coupled simulations).
  1. 1.
    Fluid dynamics solver: The solution of the discretized Navier–Stokes equations, i.e., Eq. (14) or Eq. (23), does not present significant computational difficulties, as the governing discrete operators exhibit diagonal dominance for sufficiently small time steps (due to the presence of temporal derivative terms). In the present study, we have utilized the direct method proposed by [42], as implemented in [43], for factorizing the Helmholtz operator. However, previous experience suggests that standard Krylov subspace solvers, such as BiCGStab or GMRES, can also be employed effectively, typically converging within two or three iterations. Another feasible approach involves approximating the inverse Helmholtz operator via an $N$-th order Taylor series expansion, exploiting its symmetry and positive definiteness for an appropriate choice of the time step $\Delta t$ and approximation order $N$ [44].
  2. 2.
    Structural dynamics solver: in addition to the fluid equations, the system also includes equations governing structural dynamics — specifically Eqs. (22), (24), and (25) — relevant for two-way coupled fluid-structure interaction. These equations represent initial-value ordinary differential equations (ODEs), and their numerical solution is straightforward and computationally inexpensive. Currently, we discretize them using a second-order backward finite-difference scheme, although alternative methods, such as those from the Runge-Kutta family, or even analytical solutions could equally be considered.
  3. 3.
    The pressure and force-density corrections coupled system: Ultimately, the computational bottleneck and the cornerstone of our numerical framework is the efficient solution of the regularized saddle-point linear system given by either Eq. (20) for one-way coupling or Eq. (27) for two-way coupling scenarios. Since this system appears at each iteration or time step, its efficient solution has a significant impact on the overall computational cost and stability of the algorithm.
Let us examine more closely the challenges associated with solving the saddle-point system. A common approach is to solve the full coupled system directly, as originally proposed in the IBPM by [21]. In their formulation, the authors derived a modified Poisson equation by approximating the inverse Laplacian operator via a Taylor series expansion in time. Notably, when a first-order approximation is applied, it leads to a saddle-point system with the same left-hand-side as in the present work. The authors note that given a careful choice of time step and Lagrangian points, the modified Poisson operator is symmetric and positive definite, and hence the system can be solved using the conjugate gradient method. Nevertheless, the efficiency of the iterative solution is not directly addressed, and there is no guarantee for iteration count independent on grid resolution and physical parameters. A key challenge of this approach lies in the involvement of the Laplacian operator, which inherently governs the system. The Laplacian is known for its poor conditioning due to the elliptic nature of the diffusion process it represents. It spans multiple spatial scales, resulting in a wide eigenvalue spectrum and inherently poor conditioning. This limitation becomes more pronounced at higher resolutions, as finer grids capture smaller spatial scales, leading to unbounded growth in the condition number. Despite the existence of efficient method for the solution of standard Poisson problems, such as multigrid method and specially tailored direct solvers, these methods are inefficient when applied directly to the modified Poisson equation arising in our saddle-point formulation. These considerations underscore the need for an efficient preconditioner to enable robust and rapid convergence of solutions to the saddle-point systems described by Eq. (20) or Eq. (27), as well as convergence theory. Particularly, if this preconditioner enables reduction of modified to standard Poisson problems, it will allow the efficient use of many existing methods. The development of such a preconditioner, along with both formal and numerical validation of its effectiveness, is the focus of the following sections.
@@ 4.1 The Laplacian as a preconditioner
The idea of constructing an efficient preconditioner originates from forming the primal Schur complement:
@@
and subsequently solving the decomposed system of equations, first for the pressure correction $p^{\prime}$ and then for the force-density correction $\mathbf{F}^{\prime}$:
$S_{p}p^{\prime}$ $=RHS_{p^{\prime}}+B^{T}C^{-1}RHS_{\mathbf{F}^{\prime}},$ (31)
$\mathbf{F}^{\prime}$ $=C^{-1}(Bp^{\prime}-RHS_{\mathbf{F}^{\prime}}).$ (32)
To reduce computational cost, we adopt the scalar approximation $C\approx\frac{1}{2}I_{m}$ proposed in [30], thereby avoiding the explicit inversion of $C$ by using $C^{-1}\approx 2I_{m}$. This leads to the following approximation of the primal Schur complement:
@@
Substituting this approximation into Eqs. (31) and (32) yields the simplified system:
$\tilde{S}_{p}p^{\prime}$ $=RHS_{p^{\prime}}+2B^{T}RHS_{\mathbf{F}^{\prime}},$ (34)
$\mathbf{F}^{\prime}$ $=2(Bp^{\prime}-RHS_{\mathbf{F}^{\prime}}).$ (35)
However, both $S_{p}$ and $\tilde{S}_{p}$ remain large and ill-conditioned, primarily due to the system’s dependence on the Laplacian operator. To address this, we explicitly employ the Laplacian $L$ itself as a preconditioner. In the following sections, we demonstrate that Eq. (31) can be efficiently solved using a Krylov subspace method preconditioned by $L$, leveraging the spectral equivalence between $S_{p}$ and $L$. Specifically, in Section 5, we show that the eigenvalues of the preconditioned operator $L^{-1}S_{p}$ are contained within a narrow, bounded interval that is independent of physical parameters and grid resolution. Provided that $\tilde{S}_{p}$ closely approximates $S_{p}$, similar spectral properties hold approximately for the operator $L^{-1}\tilde{S}_{p}$. This spectral equivalence leads to rapid convergence of Eq. (34) when solved with the Laplacian preconditioner, as verified numerically in Section 6. Since the pressure–force-density correction system constitutes the main computational bottleneck, this preconditioning strategy substantially improves overall computational efficiency and demonstrates favorable scalability.
@@
There is abundance of literature on block-preconditioning for saddle point matrices as appears in Eq. (20) or (27), mainly by Schur complement preconditioning. Since forming and inverting the exact Schur-complement is computationally expensive, broad research is dedicated to the search of cheap approximations for the inverse of the Schur-complement, see [45]. Our method offers an efficient approximation for the primal Schur complement for a certain class of saddle point matrices, and can be implemented either by the steps described in Section 3 or by a corresponding block lower triangular Schur complement preconditioner.
4.2 Efficient matrix-free implementation
In this subsection, we describe the current matrix-free implementation. First, recall that the operator $B^{T}$ consists of the (negative) divergence applied to the discrete Dirac delta function weights obtained from Eq. (11). A matrix-free implementation allows efficient shared-memory parallelization (implemented here with OpenMP and executed with 48 threads) and enables rapid recomputation of these operators at each time step. Second, at each step of the preconditioned Krylov iteration, only a matrix-vector product with the preconditioned operator is required. We employ left preconditioning, which is mathematically equivalent to solving Eq. (34) after multiplying both sides from the left by $L^{-1}$. However, we do not explicitly form the inverse operator $L^{-1}$, nor do we repeatedly apply a Laplacian solver to multiple columns. Instead, noting that the preconditioned system can be represented as:
@@ 4.2 Efficient matrix-free implementation
where $I_{m}$ is an identity matrix of appropriate dimension, we observe that the product of $L^{-1}\tilde{S}_{p}$ with a vector $\mathbf{v}$ can be efficiently obtained by applying the direct solver of [42] only once to the vector $B^{T}(B\mathbf{v})$. Employing Intel MKL routines [^4] We convert $B$ and $B^{T}$ into CSR format at each time step to facilitate efficient use of Intel MKL routines. allows efficient computation of the sparse matrix-vector products. Furthermore, due to the scalar approximation for $C^{-1}$, computing the second term of Eq. (34) and the entire Eq. (35) is trivial, as both involve only a single matrix-vector product. Importantly, our matrix-free implementation requires only a single application of the operator $L^{-1}$ to the vector $B^{T}(B\mathbf{v})$ per iteration or time step. Consequently, it incurs no additional computational cost compared to non-preconditioned solvers for Eqs. (20) or (27).
Finally, we discuss the factorization of the preconditioner $L$. Lynch [42] demonstrated that any matrix expressible as a sum of Kronecker products of smaller matrices with identity matrices can be factorized efficiently using eigen-decomposition of each smaller matrix. As shown therein, the resulting system can be solved with a computational complexity of $O(n^{4/3})$ for a regular 3D grid. The computational cost can be further reduced by applying the Thomas TDMA algorithm along one of the three coordinate directions. For transient problems, this approach is highly efficient, as the associated operators can be factorized once in advance and reused at every time step [43].